The cars initial velocity is 40km/h. The velocity has sustained until the 3rd hour. However,the cars velocity changes to 20km/h after 4 hours and finally stops at the 5th hour. What is the displacement of the car?
1
Expert's answer
2020-12-04T11:17:08-0500

The displacement of the car after the first 3 hours is:

"d_1 = 40km/h\times 3h = 120km"

The displacement of the car during the 4-th hour is:


"d_2 = 20km/h\times 1h = 20km"

And the displacement of the car duing the 5-th hour is zero, for the car stops. Thus, the final displacement is:


"d = d_1 + d_2 =140km"

Answer. 140 km.